# A Statistical Machine Learning Approach for Adapting Reduced-Order Models using Projected Gaussian Process

### v1

@@ 4 Numerical Examples
Three numerical experiments are performed to demonstrate the proposed method: (i) flow around a circular cylinder; (ii) heat transfer in powder bed fusion additive manufacturing; and (iii) advection-diffusion processes.
4.1 Example I: Flow around a Cylinder
We first introduce the experiment setup and then present the experiment results.
Experiment Setup. In Example I, we consider the flow passing over a cylinder in a rectangle domain. The Lattice Boltzmann method is used to simulate the flow, i.e., to generate the training data for our model [22]. In particular, the Bhatnagar–Gross–Krook (BGK) approximation to the Navier-Stokes equation is:
(32) $\dot{f}+(\bm{v}\cdot\triangledown)f=-\tau(f-\bar{f})$
where $f(\bm{x})$ be the phase-space density of fluids at location $\bm{x}$, $\bm{v}$ is the velocity, $\tau$ is the timescale of which collisions happen, and $\bar{f}$ represents the equilibrium state. In (32), the left side $(\frac{\partial}{\partial t}f+(\bm{v}\cdot\triangledown))f$ represents the streaming of particles, while the right side $-\tau(f-\bar{f})$ represents the collision. The model assumes that the fluid locally relaxes to equilibrium over a characteristic timescale $\tau$ related to the kinematic viscosity. The larger the $\tau$ is, the larger is the kinematic viscosity.
Figure 5: Simulated flow velocity at times ($\tau=0.53$)
The fluid equation (32) is solved by the Lattice Boltzmann method using the Python code in [22]. As shown in Figure 5, we consider a $400\times 100$ rectangle domain. A solid cylinder is placed at location $(100,50)$ with a radius of 13. The velocity goes from left to right with a magnitude of $v=2.3$ per time step, and 5000 time steps are simulated. The initial condition follows a Gaussian distribution with mean zero and standard deviation 0.01. Reflective boundary conditions are considered for the cylinder, the top and the bottom boundaries, while the fluid is absorbed by the right boundary. Figure 5 shows the velocity of the flow, with $\tau=0.53$, at times 1, 2000, 4000 and 5000. The Von Karman vortices are visible at times 4000 and 5000.
Figures 6 and 7 show the 1st, 4th, 7th and 10th POD modes for $\tau=0.53$ and $\tau=0.7$, respectively. We see that the POD modes change as the parameter $\tau$ changes, suggesting the necessity of adapting POD basis.
Figure 6: The 1st, 4th, 7th and 10th POD modes for $\tau=0.53$
Figure 7: The 1st, 4th, 7th and 10th POD modes for $\tau=0.7$
To evaluate the accuracy of the predicted ROM through cross validation, we generate data for 33 values of $\tau$ ranging from 0.53 to 0.7 with a step size of 0.005. Then, the data are divided into a training and a testing sets. The training set includes the data generated from the following 11 parameter settings,
$\tau^{\text{train}}=\{0.53,0.545,0.56,0.575,0.59,0.605,0.62,0.635,0.65,0.665,0.68\},$
while the data generated from the remaining 22 parameter settings are used for testing.
In particular, we let $m(\tau)$ yields a constant basepoint which is the subspace spanned by the global POD basis. The kernel $\omega$ is chosen as $\omega(\tau_{i},\tau_{j})=\xi_{1}\exp(-\xi_{2}^{-1}|\tau_{i}-\tau_{j}|)$, and $\bm{K}=\sigma_{\bm{K}}\bm{I}$. Note that, $\bm{K}_{*}(\bm{K}^{\text{train}})^{-1}$ does not depend on $\sigma_{\bm{K}}$ when $\bm{K}=\sigma_{\bm{K}}\bm{I}$, and only $\xi_{1}$ and $\xi_{2}$ are to be estimated by minimizing the cross-validation error. We obtain $\hat{\xi}_{1}\approx 1$ and $\hat{\xi}_{2}\approx 0.5$.
Experiment Results. Figure 8 shows the out-of-sample testing errors for the proposed approach, the existing POD interpolation method, and the use of global POD bases. The errors are measured by the $\ell_{2}$ Frobenious norm defined under (2). Some important observations are summarized as follows:
$\bullet$ The proposed approach yields much smaller $\ell_{2}$ error than the existing POD interpolation method from testing cases #1 to #6, as well as from #19 to #22 (i.e., towards the two boundaries of the parameter space of $\tau$). This consist observation suggests that the existing Lagrangian interpolation may have some stability issues for parameters near the boundaries of the parameter space, as investigated in [12]. In addition, as discussed in the introduction section, the proposed approach allows us to optimally estimate the model parameters based on data, while the coefficients of the Lagrangian interpolation are independent from data. Overall, the proposed approach yields smaller $\ell_{2}$ error than the existing POD interpolation method for 21 out of the 22 testing cases (except for testing case #14 where the $\ell_{2}$ errors are almost identical).
$\bullet$ The benefit of interpolating POD basis over the use of global POD bases is also demonstrated. The proposed approach yields smaller $\ell_{2}$ errors than the use of the global POD bases for 17 out of the 22 testing cases.
Figure 8: Comparison of out-of-sample testing errors measured by the Frobenious norm
Figure 9: Standard deviation associated with the ROM at time 100.
As a statistical learning approach, the proposed approach is capable of quantifying the uncertainty of the output generated from the ROM. Figure 9 shows the bootstrapping standard deviation of the output of the ROM for testing cases #1, #8, #15 and #22. Within each iteration, a subspace is randomly sampled from the distribution $\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{u}_{*},\tilde{\bm{K}}_{*}))$ obtained from (28), and an ROM is constructed on that sampled space. This procedure is repeated for 1000 times, and the standard deviation of the ROM output at time 100 is computed for each spatial location.
4.2 Experiment II: Powder Bed Fusion Additive Manufacturing
@@
Powder Bed Fusion (PBF) with metals or polymers is one of the most common 3D printing technologies in industrial additive manufacturing (AM). PBF works by jointing powdered material point by point using energy sources such as lasers or electron beams. Thermal analysis is essential for product quality control which is key to spreading laser-based AM techniques [33] and manufacturing process design [17].
@@
Experiment Setup. In this example, we focus on the transient heat transfer in the single-track L-PBF (laser PBF) process as shown in Figure 10. The vacuum chamber is preheated to a specific temperature and the spatial domain is a combination of the solid substrate and the powder layer. The transient heat transfer process is governed by a PDE:
(33) $\rho c_{p}\frac{\partial u(x,t)}{\partial t}-\nabla\cdot k\nabla u(x,t)=0,\quad(x,t)\in\Omega\times[0,t_{f}],$
(34) $(-k\nabla u(x,t))\cdot\mathbf{n}=q_{h}+q_{c}+q_{r},\quad(x,t)\in\partial\Omega\times[0,t_{f}],$
@@
(35) $q_{h}=-\frac{2\alpha\tau}{\pi\gamma^{2}}\text{exp}(-\frac{2\|x-x^{c}(t)\|^{2}}{\gamma^{2}}),\quad x\in\partial\Omega_{t},$
@@
(36) $q_{c}(x,t)=h(u(x,t)-u_{a}),\quad q_{r}(x,t)=\sigma_{s}\epsilon(u(x,t)^{4}-u_{a}^{4}),$
@@
The Finite Element method (FEM) is used to solve the governing equations. Multiplying both sides of (33) by a test function $\phi$ and integrating over $\Omega$ gives
[…]
@@
[…]
@@
[…]
@@
We generate data for 7 laser power levels (in Watt): $\{500,550,600,650,700,750,800\}$, using the Finite Element method. As an illustration, Fig.11 shows the temperature field at selected times. Since seven laser power levels are used to generate the data, a Leave-One-Out Cross-Validation (LOOCV) is performed to compare the model performance. Similar to Example 1, we let $r=10$ and $m(\tau)$ be a constant basepoint which is the subspace spanned by the global POD basis. The kernel $\omega$ is chosen as $\omega(\tau_{i},\tau_{j})=\xi_{1}\exp(-\xi_{2}^{-1}|\tau_{i}-\tau_{j}|)$, and $\bm{K}=\sigma_{\bm{K}}\bm{I}$.
@@
$\bullet$ More importantly, the proposed pGP is able to provide an explanation to such an observation. In short, this is because the proposed pGP (i) involves the use of global POD basis as its special case (due to its parameter selection capabilities), and (ii) successfully identifies that the ROM, in this example, is not sensitive to the change of parameters, and the use of the global POD basis becomes appropriate.
To elaborate, the kernel $\omega$ is chosen as $\omega(\tau_{i},\tau_{j})=\xi_{1}\exp(-\xi_{2}^{-1}|\tau_{i}-\tau_{j}|)$ in this example. When LOOCV is performed, we note that the estimated values for $\xi_{2}$ ranges between 40 to 80, effectively making $\omega(\bm{\theta}_{i},\bm{\theta}_{j})$ close to zero. In this case, (28) immediately implies that $\bm{u}_{*}=\bm{0}$ and $p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}_{*}))$. Because the global POD basis is chosen as the reference point, then, any subspace (corresponding different power levels $\tau$) on the Grassmann Manifold can be seen as i.i.d. samples from a PGD, $\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}))$. Hence, the point prediction, $\hat{p}_{*}$, is close to the subspace spanned by the global POD basis, justifying the observation in Figure 12.

### v2

@@ 4 Numerical Examples
Numerical investigations are performed to demonstrate the performance of the proposed pGP approach. In particular, we compare the prediction accuracy between the proposed method and the existing ones, illustrate the uncertainty quantification capabilities and how the pGP can help to identify important parameters.
Experiment setup and data. Consider, $\mathcal{A}({\bm{\theta}})x(t,\bm{s})=0$, an advection-diffusion process, where $x(t,\bm{s})$ is a spatio-temporal process in space $\bm{s}\in\mathbb{S}\subset\mathbb{R}^{2}$ and time $t\in[0,T]$, and $\mathcal{A}({\bm{\theta}})$ is the advection-diffusion operator parameterized by $\bm{\theta}$:
$\mathcal{A}x:=\dot{x}+\bm{\vec{v}}^{T}\triangledown x-\triangledown\cdot[\bm{D}\triangledown x]$ (32)
where $\bm{\vec{v}}$, $\bm{D}$, $\triangledown$ and $\triangledown\cdot$ represent the velocity, diffusivity, decay, gradient and divergence, respectively. In this experiment, $\mathbb{S}$ is a 2D rectangular domain.
To investigate the performance of adapting the POD basis using the proposed pGP, we first let $\bm{v}=(1,0)^{T}$ and $\bm{D}=\mathrm{diag}\{d_{1},d_{2}\}$, and generate data by letting $(d_{1},d_{2})$ take values from a 2D mesh grid with uniformly spaced $x$- and $y$-coordinates in the interval $[0.01,0.05]$ with a step size of 0.002 (more parameters will be included in subsequent investigations). Hence, a number of $21\times 21=441$ parameter settings are considered, and the process is simulated for each parameter setting.
For illustrative purposes, Figure 5 shows two simulated processes respectively for two parameter settings $[0.01,0.01]$ and $[0.05,0.05]$ of $(d_{1},d_{2})$, and Figure 6 shows the four leading POD modes of the two processes. It is seen that the POD modes vary as the parameters $d_{1}$ and $d_{2}$ change, suggesting the potential need to adapt the POD basis for parameters change.
Figure 5: Two simulated processes respectively at parameter settings $[0.01,0.01]$ (top) and $[0.05,0.05]$ (bottom) of $(d_{1},d_{2})$ .
Figure 6: The four leading POD modes obtained from the simulated processes at parameter settings $[0.01,0.01]$ (top) and $[0.05,0.05]$ (bottom) of $(d_{1},d_{2})$.
Performance metrics. To comprehensively evaluate the accuracy of the predicted POD bases, four performance metrics are adopted in our investigations:
[…]
$\bullet$ Following the $\ell_{2}$ error above, we also report the relative error so as to assess the percentage increase of the $\ell_{2}$ error from the optimal to the predicted POD basis,
$e_{R}(\bm{\theta})=\frac{e_{F}(\bm{\theta})-e^{*}(\bm{\theta})}{e^{*}(\bm{\theta})}\times 100\%$ (34)
[…]
$\bullet$ In addition, motivated by the subspace angle interpolation of POD bases, the third performance metric is obtained by computing the geometric distance, i.e., the principal angel, between the subspaces respectively spanned by the optimal POD basis and the predicted POD basis.
$e_{A}(\bm{\theta})=\left(\sum_{i=1}^{r}\text{arccos}^{2}(\sigma_{i})\right)^{\frac{1}{2}}$ (35)
[…]
$\bullet$ Finally, we evaluate the worst case scenario by computing the $\ell_{\infty}$ prediction error as follows:
[…]
where $||\cdot||_{\infty}$ is the matrix infinity norm which shows the largest error among the columns of the matrix $\bm{D}(\bm{\theta})-\hat{\bm{\Phi}}(\bm{\theta})\hat{\bm{\Phi}}(\bm{\theta})^{T}\bm{D}(\bm{\theta})$.
Numerical results. The data are divided into a training set and a testing set. The training set includes 36 parameter settings $\bm{\theta}=(d_{1},d_{2})\in\Theta^{\text{train}}$ from a coarse 2D mesh grid with uniformly spaced $x$- and $y$-coordinates in the interval $[0.01,0.05]$ with a larger step size of 0.008. Data generated from the remaining 405 parameter settings, $\bm{\theta}\in\Theta^{\text{test}}$, are reserved for testing.
Figure 7 shows the four types errors, $e_{F}$, $e_{R}$, $e_{A}$ and $e_{I}$, for all 405 testing cases while letting the reduced dimension to be $r=5$. Here, three different approaches are used, including the POD interpolation (using a Gaussian radial kernel), the global POD approach, and the proposed pGP approach. It is noted that, the proposed pGP approach yields the best performance in all four performance metrics. In particular,
$\bullet$ In terms of $e_{F}$ and $e_{R}$, the proposed pGP outperforms the interpolation approach and the global POD approach respectively in 365 and 395 out of the 405 testing cases.
$\bullet$ In terms of $e_{A}$, the proposed pGP outperforms the interpolation approach and the global POD approach in 351 and 385 out of the 405 testing cases, respectively.
$\bullet$ In terms of $e_{I}$, the proposed pGP outperforms the interpolation approach and the global POD approach in 372 and 349 out of the 405 testing cases, respectively.
Figure 7: Testing errors, $e_{F}$, $e_{R}$, $e_{A}$ and $e_{I}$, for all 405 testing cases ($r=5$)
Figure 8: Testing errors, $e_{F}$, $e_{R}$, $e_{A}$ and $e_{I}$, for all 405 testing cases ($r=10$)
Next, we increase the reduced dimension to $r=10$, and the updated comparison results are shown in Figure 8. The proposed pGP still yields the best overall performance. We also note that, when the dimension of the reduced-order basis increases, all three approaches yield smaller errors in terms of $e_{F}$, $e_{A}$ and $e_{I}$. However, in terms of the relative error $e_{R}$, the errors seem to become larger. This is not surprising as further investigation shows that the first 10 optimal POD modes (obtained from performing SVD directly on the testing snapshot data) can already capture about 97% of the total energy, making the denominator $e^{*}$ in (34) very small.
Figure 9: The proposed pGP includes the global POD approach as its special case by yielding almost identical $e_{F}$ errors for all 405 testing cases when the dimension of the reduced-order basis is increased to $r=20$.
Relationship between the proposed and global POD approaches. Following the discussions above, one may wonder what if the dimension of the reduced-order basis is further increased. Note that, an important property of the proposed pGP is that it includes the use of global POD basis as its special case. As already discussed in Sec.3.4, consider the squared Exponential kernel, $\omega(\bm{\theta}_{i},\bm{\theta}_{j})=\xi_{1}^{2}\delta_{\{\bm{\theta}_{i}=\bm{\theta}_{j}\}}+\xi_{2}^{2}\exp\{-(\bm{\theta}_{i}-\bm{\theta}_{j})^{T}\bm{K}_{\omega}(\bm{\theta}_{i}-\bm{\theta}_{j})/2\}$, where $\bm{K}_{\omega}=\text{diag}\{\xi_{3}^{-2},\xi_{4}^{-2}\}$. If the estimated $\xi_{1}$, $\xi_{3}$ and $\xi_{4}$ are all very small, then, $\omega(\bm{\theta}_{i},\bm{\theta}_{j})$ is close to zero and it is implied by (28) that $\bm{u}_{*}=\bm{0}$ and $)p_{*}|\bm{p}^{\text{train}}\sim\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}_{*}$)). Because the global POD basis is chosen as the reference point, any subspace (corresponding to different $\bm{\theta}\in\Theta^{\text{test}}$) on the Grassmann manifold can be essentially seen as the i.i.d. samples from a pGD, $)\mathcal{P}_{m_{*}}(\mathcal{MN}(\bm{0},\tilde{\bm{K}}$)). Hence, the point prediction, $\hat{p}_{*}$, is close to the subspace spanned by the global POD basis. This theoretical result is illustrated in Figure 9, which shows that the proposed pGP yields almost identical $e_{F}$ errors as the global POD approach does for all 405 testing cases when the dimension of the reduced-order basis is increased to $r=20$.
In the Appendix, we provide one more example that shows how the proposed pGP includes the use of global POD as its special case.
Figure 10: Bootstrapping standard deviation of the predicted subspace over the parameter space: the uncertainty becomes larger for parameter settings away from training points indicated by “++”.
[…]
To illustrate the output generated by the procedure above, Figure 10 shows the bootstrapping standard deviation of the predicted subspace at each testing parameter setting over the parameter space (here, $\tilde{n}=1000$). In this figure, the small cross “++” indicates the training points where there is no uncertainty. A clear pattern is observed: the uncertainty is the lower for parameter settings closer to the training points, while the uncertainty becomes larger for parameter settings away from training points. In general, the uncertainty is the largest halfway between existing training points. Such a pattern is consistent with our intuition.
Parameter Selection Capabilities. As discussed in Sec.3.4, it is often the case that not all parameters are equally important in terms of their predictive power for POD basis modes. In other words, the POD modes can be sensitive to some parameters while independent of the others. Note that, this idea may seem to share some similarities with the concept of active subspace (Constantine et al.,, 2014). The difference between the concept of active subspace and parameter selection discussed here is that, the former identifies a low-dimensional space of model input which contains the detected directions of the strongest variability in model output, while the latter refers to the process of identifying the parameters which have the strongest predictive power of the POD modes in a statistical model (for a specific problem and given a specific training dataset).
To demonstrate the parameter selection capabilities, we expand the parameter space and consider the change of both the velocity $\bm{v}=(v_{1},v_{2})$ and diffusivity $\bm{D}=\text{diag}\{d_{1},d_{2}\}$. In particular, the training data are simulated for $(v_{1},v_{2})$ values from a 2D mesh grid with uniformly spaced x- and y-coordinates in the interval [-0.1, 0.1] with a step size of 0.05, and for $(d_{1},d_{2})$ values from a 2D mesh grid with uniformly spaced x- and y-coordinates in the interval [0.01, 0.05] with a step size of 0.05. Before the pGP model is trained, the values of each of the four parameters are standardized between 0 an 1, and the standardized parameter space is $0,1]^{4}$.
In Rasmussen and Williams, (2005), the squared Exponential kernel is considered when introducing the parameter selection problem for GP. In our experiment, we adopt the same type of kernel as follows with a minor modification:
$\omega(\bm{\theta}_{i},\bm{\theta}_{j})=\xi_{1}^{2}\delta_{\{\bm{\theta}_{i}=\bm{\theta}_{j}\}}+\xi_{2}^{2}\exp\left\{-\frac{(\bm{\theta}_{i}-\bm{\theta}_{j})^{T}\bm{K}_{\omega}(\bm{\theta}_{i}-\bm{\theta}_{j})}{2}\right\}.$ (37)
where $\bm{\theta}=(v_{1},v_{2},d_{1},d_{2})$, $\bm{K}_{\omega}=\text{diag}\{\xi_{3}^{-2},\xi_{3}^{-2},\xi_{4}^{-2},\xi_{4}^{-2}\}$, and $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ are the parameters in the kernel function to be estimated from data. Note that, $\xi_{3}$ and $\xi_{4}$ are known as the characteristic length-scales for velocity and diffusivity parameters. Following the idea of automatic relevance determination, a smaller length-scale indicates a stronger effect of a parameter on the prediction (Neal,, 2012). One may also note that we let $v_{1}$ and $v_{2}$ share the same length-scale $\xi_{3}$, and let $d_{1}$ and $d_{2}$ share the same length-scale $\xi_{4}$. This set up is appropriate considering the parameter mesh grids used to generate the training data, and there is no strong reason to justify why $v_{1}$ and $v_{2}$ (or, $d_{1}$ and $d_{2}$) should have very different length-scales.
Figure 11: Growth of the estimated length-scales $\hat{\xi}_{3}$ and $\hat{\xi}_{4}$ as the tuning parameter $\gamma$ increases.
The parameters $(\xi_{1},\xi_{2},\xi_{3},\xi_{4})$ are estimated by the maximum likelihood method (Rasmussen and Williams,, 2005). In particular, to demonstrate the parameter selection capabilities, an $\ell_{2}$ regularization is imposed on $\xi_{3}$ and $\xi_{4}$, i.e., $\gamma(\xi_{3}^{-2}+\xi_{4}^{-2})$ for some tuning parameter $\gamma\geq 0$. Note that, (i) the regularization is only imposed on the length-scales $\xi_{3}$ and $\xi_{4}$, and (ii) as $\gamma$ increases, the regularization tends to make both $\xi_{3}$ and $\xi_{4}$ larger, i.e., to make the parameters (features) less influential in the regression model. Hence, the length-scale associated with the parameter that has less predictive power is expected to grow faster, helping us identify the important parameters.
Figure 11 shows the estimated length-scales $\hat{\xi}_{3}$ and $\hat{\xi}_{4}$ for five different values of $\gamma$, i.e., 0, 500, 1000, 1500 and 2000. It is seen that, (i) when $\gamma=0$ (i.e., no regularization is added), the estimated values of $\xi_{3}$ and $\xi_{4}$ are respectively $\hat{\xi}_{3}=0.249$ (length-scale associated with the velocity parameter) and $\hat{\xi}_{4}=0.208$ (length-scale associated with the diffusion parameter). Because $\hat{\xi}_{3}$ is slightly larger than $\hat{\xi}_{4}$, indicating that the diffusion parameter may have a slightly higher predictive power than the velocity parameter; (ii) when $\gamma$ increases, both $\hat{\xi}_{3}$ and $\hat{\xi}_{4}$ become larger as expected; (iii) when $\gamma=2000$, the estimated length scale $\hat{\xi}_{3}=1.199$ is about 2.6 times higher than $\hat{\xi}_{4}=0.456$, clearly indicating that the diffusion parameter has a higher predictive power than the velocity parameter over the parameter space considered in this experiment. The observations above demonstrate the parameter selection capabilities of the proposed pGP approach when constructing a predictive model for the POD modes.
5 Conclusions
This work successfully demonstrated the great potential of introducing machine learning to the field of robust reduced-order modeling. A supervised learning problem, based on the proposed pGP regression, has been developed to predict the optimal POD basis. The proposed approach leverages an injective mapping from the parameter space to the Grassmann manifold that contains the optimal vector subspaces. The condition for such a mapping to be injective has been obtained, and sufficient technical details were presented for predicting the POD basis using the proposed pGP regression. Numerical experiments showed that the proposed pGP approach not only improves the accuracy of ROM against parameter change, but also enables uncertainty quantification and parameter selection, potentially reducing the dimension of the parametric space when adapting the POD basis. We would also like to point out that there exists a critical link between the proposed pGP and statistical experimental design/reinforcement learning that enable sequential exploration of a system’s behavior in a reduced parameter space. This link is expected to create important future research trajectories, potentially attracting more researchers into this field and making tangible contributions to Scientific Machine Learning research and its applications. Code and data are available on GitHub: https://github.com/dnncode/pGP.
Appendix A Heat process in powder bed fusion
In the Appendix, we provide one more example that shows how the proposed pGP includes the use of global POD basis as its special case, and yields more accurate predictions of the POD basis than interpolation.
@@ Appendix A Heat process in powder bed fusion
Experiment setup. We focus on the transient heat transfer in the single-track L-PBF (laser Powder Bed Fusion) process as shown in Figure 12. PBF with metals or polymers is one of the most common 3D printing technologies in industrial additive manufacturing (AM). PBF works by jointing powdered material point by point using energy sources such as lasers or electron beams. Thermal analysis is essential for product quality control which is key to spreading laser-based AM techniques (Yan et al.,, 2018) and manufacturing process design (Liao et al.,, 2022).
The vacuum chamber is preheated to a specific temperature and the spatial domain is a combination of the solid substrate and the powder layer. The transient heat transfer process is governed by a PDE:
$\rho c_{p}\frac{\partial u(x,t)}{\partial t}-\nabla\cdot k\nabla u(x,t)=0,\quad(x,t)\in\Omega\times[0,t_{f}],$ (38)
$(-k\nabla u(x,t))\cdot\mathbf{n}=q_{h}+q_{c}+q_{r},\quad(x,t)\in\partial\Omega\times[0,t_{f}],$ (39)
@@
$q_{h}=-\frac{2\alpha\tau}{\pi\gamma^{2}}\text{exp}(-\frac{2\|x-x^{c}(t)\|^{2}}{\gamma^{2}}),\quad x\in\partial\Omega_{t},$ (40)
@@
$q_{c}(x,t)=h(u(x,t)-u_{a}),\quad q_{r}(x,t)=\sigma_{s}\epsilon(u(x,t)^{4}-u_{a}^{4}),$ (41)
@@
The finite element method is used to solve the governing equations. Multiplying both sides of (38) by a test function $\phi$ and integrating over $\Omega$ gives
$\int_{\Omega}\rho c_{p}\frac{\partial u(x,t)}{\partial t}\phi dV-\int_{\Omega}\nabla\cdot k\nabla u(x,t)\phi dV=0,\quad(x,t)\in\Omega\times[0,t_{f}],$ (42)
@@
$\int_{\Omega}\rho c_{p}\frac{\partial u(x,t)}{\partial t}\phi dV+\int_{\Omega}k\nabla u(x,t)\cdot\nabla\phi dV+\int_{\partial\Omega}(-k\nabla u(x,t))\cdot\mathbf{n}\phi dS=0,$ (43)
@@
$\rho c_{p}\sum_{i=1}^{N}\frac{\partial u_{i}(t)}{\partial t}\int_{\Omega}\psi_{i}\psi_{j}dV+k\sum_{i=1}^{N}u_{i}(t)\int_{\Omega}\nabla\psi_{i}\nabla\psi_{j}dV+\sum_{i=1}^{N}\int_{\partial\Omega}(-k\nabla u_{i}(t)\psi_{i})\cdot\mathbf{n}\psi_{j}dS=0$ (44)
@@
We generate data for 7 laser power levels (in Watt): $\{500,550,600,650,700,750,800\}$, using the finite element method. As an illustration, Figure 13 shows the temperature field at selected times. Since seven laser power levels are used to generate the data, the Leave-One-Out Cross-Validation (LOOCV) is performed to compare the model performance. We let $r=10$ and $m(\tau)$ be a constant basepoint which is the subspace spanned by the global POD basis. The kernel $\omega$ is chosen as $\omega(\tau_{i},\tau_{j})=\xi_{1}\exp(-\xi_{2}^{-1}|\tau_{i}-\tau_{j}|)$, and $\bm{K}=\sigma_{\bm{K}}\bm{I}$.
@@
$\bullet$ More importantly, the proposed pGP is able to provide an explanation to such an observation. In short, this is because the proposed pGP (i) involves the use of global POD basis as its special case (due to its parameter selection capabilities), and (ii) successfully identifies that the POD basis, in this example, is not sensitive to the change of parameters, and the use of the global POD basis becomes appropriate.